# Sparsity May Be All You Need: Sparse Random Parameter Adaptation

### v2

@@ 7.2 Which parameters should SpaRTA target?
Now, we investigate whether it is better to concentrate the selection of sparse, trainable parameters on a few parameters or equally across all. Specifically, given a budget for the number of parameters that can be trained, which parameters should we target for sparsification (while freezing the remaining) to achieve best task performance when adapting the sparsified parameters within the targeted types?
We conduct an ablation study to answer this question. The Gemma 2B model is adapted to the SST-2 task using our SpaRTA approach targeting different combinations of parameters. We set a budget of 1.25M trainable parameters within the Gemma 2B model so that sparsity $s=99.95\%$.
@@
Note that concentrating our sparsification over targeted parameters, although proven beneficial in terms of single-task adaption performance, can decrease performance when merging since it increases the chance of parameter interference.
@@ 7.3 GLUE and BoolQ
Model: Mistral-7B-Instruct-v0.3 QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ method targets loss acc. loss acc. loss acc. loss acc. loss acc. loss acc. loss mcc loss acc. avg. PT (zero-shot) 2.38 23.4 2.54 33.6 0.66 66.3 0.85 57.5 2.42 35.5 0.70 70.0 0.72 25.2 0.39 85.2 49.6 SpaRTA (5%) ALL 0.11 96.0 0.26 89.8 0.12 95.9 0.22 90.7 0.23 91.6 0.32 87.3 0.36 62.6 0.24 90.8 88.1 SpaRTA (0.5%) 0.11 95.8 0.30 90.5 0.12 96.0 0.23 90.4 0.24 91.2 0.32 88.0 0.35 67.9 0.24 90.5 88.8 SpaRTA (0.048%) 0.12 95.7 0.31 89.2 0.12 96.1 0.23 90.5 0.24 91.1 0.31 88.2 0.34 67.3 0.26 89.5 88.4 LoRA (0.048%) Q,V 0.12 95.6 0.24 91.2 0.12 96.0 0.23 90.2 0.24 91.0 0.29 89.3 0.32 69.9 0.24 90.5 89.2 DoRA (0.048%) 0.13 95.0 0.27 90.9 0.11 96.4 0.23 90.2 0.25 90.9 0.28 88.7 0.32 68.2 0.24 90.5 88.8 SpaRTA (0.048%) 0.12 95.6 0.31 88.3 0.13 95.6 0.23 90.1 0.25 90.9 0.30 88.6 0.35 65.2 0.26 89.7 88.0 SpaRTA (0.048%) O,V 0.11 95.9 0.31 88.4 0.14 95.8 0.23 90.4 0.24 91.2 0.30 88.4 0.34 64.9 0.26 89.7 88.1
Table 6: Test loss and accuracy of the mistral-7b-it model adapted to GLUE and BoolQ datasets with different fine-tuning methods. For training details see Appendix E.2.
@@
Table 7: Test loss and accuracy of models adapted to the MMLU dataset with SpaRTA and LoRA. We also report zero-shot performance of instruction following PT models. For training details see Table 17 in the Appendix.
We now focus on comparing SpaRTA and LoRA methods using the 7 NLU task in GLUE tasks as well as BoolQ. Table 5 presents the results for adapting gemma-2b-it. An equivalent set of results for the gemma-2b base model is given in Table 10, see Appendix D. Adaptation results (rows) are grouped by which type of parameters are targeted, with SpaRTA further ordered in descending order of density. SpaRTA with 0.037% density has approximately the same number of trainable parameters as LoRA, for which a rank $r=8$ was used.
@@
A similar set of results on GLUE and BoolQ is provided for mistral-7b-it in Table 6. Again, results for the mistral-7b base model can be found in Table 11 in Appendix D. Again, SpaRTA exhibits comparable performance to LoRA and DoRA for a low density of 0.0048%, where all methods have comparable numbers of trainable parameters. And once again, with SpaRTA targeting Wv and Wo typically outperforming the targeting of Wq and Wv.
@@ 7.4 MMLU
We have chosen MMLU because it is known for being a very challenging NLU task. Table 7 compares SpaRTA against LoRA on adapting each of our models to the MMLU multiple choice question answering task, where each model must predict the correct answer to a set of questions. Solving this task requires a high level of world knowledge and problem solving skills. In this experiment, we restrict both methods to use approximately the same number of trainable parameters. See Table 17 in the Appendix for training details. Here again, the results show that SpaRTA is competitive with LoRA, providing similar performance.
@@ 7.5 Remarks
Our results establish that SpaRTA is a viable adaptation technique that can be competitive with LoRA, especially for larger LMs. These results demonstrate that a simple sparsifying scheme can offer a valid adaptation technique. This opens the possibility for further investigating sparse adaptation as the low rank approximation of a model’s parameter changes is not the only mechanism to provide a performant adaptation method. This indicates that what matters is not necessarily the adapter structure used in PEFT techniques, but rather the number of trainable parameters, that is relevant to the adaptation task at hand.
@@ Appendix A Ranks of differences in fine-tuned weight matrices
We compute all delta weight matrices for two well-known instruction-following fine-tuned models: gemma-2b-it and mistral-7b-it, see Section 6.2. As we can see in Tables 8 and 9, the feedforward (MLP) delta $\Delta$-matrices associated with each layer of the transformer network are all full rank [^5] We compute the rank of a matrix as the number of singular values that are greater than zero, given a specified tolerance, using torch.linalg.matrix_rank. We did some sensitivity analysis to determine such tolerance. in both models. This is also the case for all self attention key (k) and value (v) projection matrices. However, the self attention query (q) and output (o) $\Delta$-matrices all show relatively small rank deficiencies: between $9$ and $2$ for the query projection matrices and between $46$ and $3$ for the output projection matrices, out of a potential maximum rank of $2,048$ for the gemma-2b-it model; and between $1,788$ and $28$ for the query projection matrices and between $246$ and $10$ for the output projection matrices, out of a potential maximum rank of $4,096$ for the mistral-7b-it model. The token embedding $\Delta$-matrices of both the gemma-2b-it and mistral-7b-it models are full rank. Of the two models, mistral-7b-it is the only one that does not tie its token embeddings’ weights to its head; thus the mistral-7b-it’s model head (untied lm head) is updated independently during fine-tuning, and the resulting delta change of this matrix is also full rank. Basically, we observe that the fine-tuning changes in the weight matrices of these well known models are all full rank, with the only exception being the changes in the query (q) and output (o) projection matrices that show small rank deficiencies.
@@ Appendix C Training details
With SpaRTA, we froze the token embeddings layer, made fully trainable the classification head, and randomly chose a sparse proportion of (scalar) parameters to be trained in all the other layers of a model. We demonstrated our method with varying density levels. The total average number of trainable parameters in each case was approximately: $99$M (5% density), $10$M (0.5%) and $800$k (0.037%) for the (base and instruct) Gemma 2B models; and $349$M (5% density), $35$M (0.5%) and $3.5$M (0.048%) for the (base and instruct) Mistral 7B models.
For LoRA, we factorized the changes in the query and value self-attention projection weight matrices with rank $r=8$ decomposition matrices, which were optimized while keeping all other model parameters frozen. The number of new trainable parameters introduced by the LoRA approach in the Gemma 2B and Mistral 7B models were approximately $925$k and $3.5$M, respectively. Also, we set $\alpha=16$ to scale the LoRA adapters.
We observed overfitting when training with Full parameter FT: the development loss started deteriorating after a few epochs (e.g., approximately $2$ for SST-2) while the training loss went quickly to zero. We used a combination of early stopping, dropout and weight decay to deal with overfitting. We noticed that SpaRTA is a natural regularizer: increasing sparsity resulted in less overfitting, to a point in which there was no more overfitting (e.g., this was achieved at a sparsity $s>99\%$ for all models under consideration). In general, overfitting was more noticeable with the Mistral 7B models; as expected since they are larger than the Gemma 2B models. We also noticed that larger models require higher sparsity levels to eliminate overfitting given the same training data. On the other hand, Head adapation always resulted in significant underfitting, making this fine-tuning method always underperform.
We also observed that SpaRTA achieves a speedup over Full FT for sparsity levels greater than 90%. Thus, at some sparsity level between 80% and 90%, computing gradients on a smaller amount of trainable parameters starts to offset the extra compute of adding and subtracting the deltas before and after the forward pass respectively. These extra operations require slicing the original parameter tensors over the randomly selected indices.
@@ Appendix D Additional Experiment Results
Tables 10 and Table 11 present the complete set of results on GLUE and BoolQ for the PT base models gemma-2b and mistral-7b, respectively. Tables 12 and 13 present standard errors from corresponding experiments.
@@
Model: Mistral-7B-v0.3 QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ method Targets loss acc. loss acc. loss acc. loss acc. loss acc. loss acc. loss mcc loss acc. SpaRTA (5%) ALL 0.11 95.9 0.36 85.0 0.11 96.8 0.38 82.1 0.30 89.2 0.33 86.7 0.35 64.6 0.25 90.1 SpaRTA (0.5%) 0.12 95.8 0.44 84.4 0.11 96.7 0.22 90.6 0.26 90.4 0.32 87.0 0.35 66.8 0.26 89.9 SpaRTA (0.048%) 0.13 95.3 0.41 83.6 0.11 96.7 0.23 90.1 0.40 86.7 0.34 86.6 0.36 64.6 0.26 89.8 SpaRTA (0.048%) Q,V 0.14 94.5 0.37 84.0 0.11 96.6 0.24 89.7 0.33 88.6 0.36 85.5 0.36 65.2 0.28 89.3 LoRA (0.048%) 0.12 95.4 0.39 85.4 0.12 96.5 0.23 90.1 0.28 89.7 0.32 87.0 0.38 64.5 0.27 89.6 DoRA (0.048%) 0.12 95.6 0.41 85.3 0.12 96.8 0.24 90.1 0.28 89.7 0.33 86.4 0.34 65.8 0.26 89.9 SpaRTA (0.048%) O,V 0.15 94.5 0.41 85.6 0.21 91.7 0.24 89.6 0.33 87.8 0.35 86.4 0.39 59.8 0.29 89.3
@@
Model: gemma-2b-it QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ method targets loss acc. loss acc. loss acc. loss acc. loss acc. loss acc. loss mcc loss acc. PT (zero-shot) 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 SpaRTA (5%) ALL 0.00 0.14 0.03 1.54 0.01 0.51 0.00 0.21 0.01 0.26 0.01 0.48 0.00 0.88 0.01 0.54 SpaRTA (0.5%) 0.00 0.16 0.03 0.64 0.01 0.46 0.00 0.17 0.00 0.13 0.01 0.07 0.01 0.75 0.01 0.25 SpaRTA (0.037%) 0.00 0.12 0.01 0.55 0.00 0.17 0.00 0.18 0.00 0.14 0.00 0.45 0.01 0.44 0.00 0.09 LoRA (0.037%) Q,V 0.00 0.11 0.01 1.10 0.00 0.35 0.00 0.03 0.00 0.11 0.00 0.22 0.02 1.40 0.01 0.12 DoRA (0.037%) 0.00 0.11 0.01 0.55 0.00 0.20 0.00 0.03 0.00 0.09 0.00 0.32 0.01 1.31 0.01 0.29 SpaRTA (0.037%) 0.00 0.08 0.00 1.04 0.01 0.11 0.00 0.00 0.00 0.26 0.00 0.22 0.02 2.83 0.00 0.41 SpaRTA (0.037%) O,V 0.00 0.13 0.01 0.64 0.00 0.27 0.00 0.16 0.00 0.17 0.02 0.22 0.01 4.36 0.00 0.22
@@
Model: Mistral-7B-Instruct-v0.3 QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ method targets loss acc. loss acc. loss acc. loss acc. loss acc. loss acc. loss mcc loss acc. PT (zero-shot) 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 SpaRTA (5%) ALL 0.00 0.09 0.01 0.24 0.01 0.26 0.00 0.03 0.00 0.08 0.01 0.47 0.01 1.91 0.00 0.16 SpaRTA (0.5%) 0.00 0.05 0.01 1.07 0.01 0.18 0.00 0.10 0.00 0.09 0.02 0.25 0.01 1.28 0.00 0.09 SpaRTA (0.048%) 0.00 0.10 0.03 0.55 0.00 0.20 0.00 0.03 0.00 0.11 0.01 0.32 0.02 1.87 0.00 0.20 LoRA (0.048%) Q,V 0.00 0.16 0.01 0.60 0.01 0.44 0.00 0.06 0.00 0.13 0.01 0.33 0.01 1.16 0.00 0.17 DoRA (0.048%) 0.02 0.76 0.02 0.87 0.00 0.17 0.00 0.03 0.00 0.07 0.00 0.17 0.01 1.22 0.00 0.16 SpaRTA (0.048%) 0.00 0.06 0.01 0.73 0.00 0.23 0.00 0.04 0.00 0.05 0.01 0.43 0.01 1.17 0.00 0.27 SpaRTA (0.048%) O,V 0.00 0.02 0.01 0.42 0.01 0.34 0.00 0.07 0.00 0.13 0.02 0.24 0.01 0.97 0.00 0.23

### v3

@@ 7.2 Which Parameters Should SpaRTA Target?
Now, we investigate whether it is better to concentrate the selection of sparse, trainable parameters on a few model layers or equally across all. Specifically, given a budget for the number of parameters that can be trained, which type of parameters should we target for sparsification (while freezing the remaining) to achieve best task performance when adapting the sparsified parameters within the targeted types?
We conduct an ablation study to answer this question. The gemma-2b model is adapted to the SST-2 task using our SpaRTA approach targeting different combinations of parameter types. We set a budget of 1.25M trainable parameters within the gemma-2b model so that sparsity is $s=99.95\%$.
@@
Corresponding results confirming, as done in Hu et al. (2022), that targeting Wo and Wv is suboptimal for LoRA can be found in Appendix D. Note that concentrating our sparsification over targeted parameter tensors, although proven beneficial in terms of single-task adaption performance, can decrease performance when merging since it increases the chance of parameter interference.
@@ 7.3 GLUE and BoolQ
Model: mistral-7b-it QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ method targets loss acc. loss acc. loss acc. loss acc. loss acc. loss acc. loss mcc loss acc. avg. PT (zero-shot) 1.38 76.6 1.54 66.4 0.66 66.3 0.85 57.5 2.42 35.5 0.70 70.0 0.72 25.2 0.39 85.2 49.6 SpaRTA (5%) ALL 0.11 96.0 0.26 89.8 0.12 95.9 0.22 90.7 0.23 91.6 0.32 87.3 0.36 62.6 0.24 90.8 88.1 SpaRTA (0.5%) 0.11 95.8 0.30 90.5 0.12 96.0 0.23 90.4 0.24 91.2 0.32 88.0 0.35 67.9 0.24 90.5 88.8 SpaRTA (0.048%) 0.12 95.7 0.31 89.2 0.12 96.1 0.23 90.5 0.24 91.1 0.31 88.2 0.34 67.3 0.26 89.5 88.4 LoRA (0.048%) Q,V 0.12 95.6 0.24 91.2 0.12 96.0 0.23 90.2 0.24 91.0 0.29 89.3 0.32 69.9 0.24 90.5 89.2 DoRA (0.048%) 0.13 95.0 0.27 90.9 0.11 96.4 0.23 90.2 0.25 90.9 0.28 88.7 0.32 68.2 0.24 90.5 88.8 SpaRTA (0.048%) 0.12 95.6 0.31 88.3 0.13 95.6 0.23 90.1 0.25 90.9 0.30 88.6 0.35 65.2 0.26 89.7 88.0 SpaRTA (0.048%) O,V 0.11 95.9 0.31 88.4 0.14 95.8 0.23 90.4 0.24 91.2 0.30 88.4 0.34 64.9 0.26 89.7 88.1
Table 6: Test loss and accuracy of the mistral-7b-it model adapted to GLUE and BoolQ datasets with different fine-tuning methods. Standard errors can be found in Table 16, see Appendix D. For training details see Appendix E.2.
method targets gemma-2b-it mistral-7b-it gemma-2b mistral-7b LoRA Q, V 83.6 89.2 81.7 87.3 SpaRTA Q, V 82.8 88.0 78.3 86.7 SpaRTA O, V 84.0 88.1 83.2 85.6
Table 7: Average accuracy (%) per adapted model. Averages are taken across 8 tasks (7 GLUE and BoolQ) based on Tables 5, 6, 13, and 14. SpaRTA is shown to be competitive with LoRA, particularly when targeting Wo (O) and Wv (V) self-attention matrices, which were previously shown to be optimal for SpaRTA.
method targets gemma-2b-it mistral-7b-it gemma-2b mistral-7b LoRA Q, V 37.58 62.5 37.5 50 SpaRTA Q, V 0 0 0 31.3 SpaRTA O, V 62.5 37.5 62.5 18.8
Table 8: Win rates (%) per adapted model, computed across 8 tasks (7 GLUE tasks and BoolQ) based on Tables 5, 6, 13, and 14. SpaRTA is shown to be competitive with LoRA, particularly when targeting Wo (O) and Wv (V) self-attention matrices, which were previously shown to be optimal for SpaRTA.
@@
Table 9: Test loss and accuracy of models adapted to the MMLU dataset with SpaRTA and LoRA. We also report zero-shot performance of instruction following PT models. For training details see Table 20 in Appendix E.
We now focus on comparing SpaRTA and LoRA methods using the 7 NLU tasks in the GLUE benchmark as well as BoolQ. Table 5 presents the results for adapting gemma-2b-it. An equivalent set of results for the gemma-2b model is given in Table 13, see Appendix D. Adaptation results (rows) are grouped by which type of parameters are targeted, with SpaRTA further ordered in descending order of density. SpaRTA with 0.037% density has approximately the same number of trainable parameters as LoRA, for which a rank $r=8$ was used.
@@
A similar set of results on GLUE and BoolQ is provided for mistral-7b-it in Table 6. Again, results for the mistral-7b model can be found in Table 14 in Appendix D. SpaRTA exhibits comparable performance to LoRA and DoRA for a low density of 0.0048%, where all methods have comparable numbers of trainable parameters. And once again, with SpaRTA targeting Wv and Wo typically outperforming the targeting of Wq and Wv.
To further illustrate the competitiveness of SpaRTA with LoRA, Table 7 summarizes accuracy results across the 8 datasets while Table 8 offers another perspective by considering win-rates among the competing methods for each model. Note that LoRA exhibits best performance overall when targeting the Wq and Wv self-attention matrices as mentioned in the original LoRA paper Hu et al. (2022), and hence we target these parameters when using LoRA. While we have already explored optimal targets for SpaRTA, we include results targeting Wq and Wv for completeness. Across both tables, SpaRTA (Wo, Wv) outperforms SpaRTA (Wq, Wv) in six out of the eight scenarios. LoRA and SpaRTA are equally performant on average across the two tables, demonstrating the claim that SpaRTA is competitive with LoRA.
@@ 7.4 MMLU
We have chosen MMLU because it is known for being a very challenging NLU task. Table 9 compares SpaRTA against LoRA on adapting each of our models to the MMLU multiple-choice question answering task, where each model must predict the correct answer to a set of questions. Solving this task requires a high level of world knowledge and problem solving skills. In this experiment, we restrict both methods to use approximately the same number of trainable parameters. See Table 20 in the Appendix for training details. Here again, the results show that SpaRTA is competitive with LoRA, providing similar performance.
@@ 7.5 Remarks
Our results establish that SpaRTA is a viable adaptation technique that can be competitive with LoRA, especially for larger LMs. These results demonstrate that a simple sparsifying scheme can offer a valid adaptation technique. This opens the possibility for further investigating sparse adaptation as the low-rank approximation of a model’s parameter changes is not the only mechanism to provide a performant adaptation method. This indicates that what matters is not necessarily the adapter structure used in PEFT techniques, but rather the number of trainable parameters, that is relevant to the adaptation task at hand.
@@ Appendix A Ranks of Differences in Fine-Tuned Weight Matrices
We compute all delta weight matrices for two well-known instruction-following fine-tuned models: gemma-2b-it and mistral-7b-it, see Section 6.2. As we can see in Tables 10 and 11, the feedforward (MLP) delta $\Delta$-matrices associated with each layer of the transformer network are all full rank [^4] We compute the rank of a matrix as the number of singular values that are greater than zero, given a specified tolerance, using torch.linalg.matrix_rank. We did some sensitivity analysis to determine such tolerance. in both models. This is also the case for all self attention key (k) and value (v) projection matrices. However, the self attention query (q) and output (o) $\Delta$-matrices all show relatively small rank deficiencies: between $9$ and $2$ for the query projection matrices and between $46$ and $3$ for the output projection matrices, out of a potential maximum rank of $2,048$ for the gemma-2b-it model; and between $1,788$ and $28$ for the query projection matrices and between $246$ and $10$ for the output projection matrices, out of a potential maximum rank of $4,096$ for the mistral-7b-it model. The token embedding $\Delta$-matrices of both the gemma-2b-it and mistral-7b-it models are full rank. Of the two models, mistral-7b-it is the only one that does not tie its token embeddings’ weights to its head; thus the mistral-7b-it’s model head (untied lm head) is updated independently during fine-tuning, and the resulting delta change of this matrix is also full rank. Basically, we observe that the fine-tuning changes in the weight matrices of these well known models are all full rank, with the only exception being the changes in the query (q) and output (o) projection matrices that show small rank deficiencies.
@@ Appendix C Training Details
With SpaRTA, we froze the token embeddings layer, made fully trainable the classification head, and randomly chose a sparse proportion of (scalar) parameters to be trained in all the other layers of a model. We demonstrated our method with varying density levels. The total average number of trainable parameters in each case was approximately: 9999M (5% density), 1010M (0.5%) and 800800k (0.037%) for the (base and instruct) Gemma 2B models; and 349349M (5% density), 3535M (0.5%) and 3.53.5M (0.048%) for the (base and instruct) Mistral 7B models.
For LoRA, we factorized the changes in the query and value self-attention projection weight matrices with rank $r=8$ decomposition matrices, which were optimized while keeping all other model parameters frozen. The number of new trainable parameters introduced by the LoRA approach in the Gemma 2B and Mistral 7B models were approximately 925925k and 3.53.5M, respectively. Also, we set $\alpha=16$ to scale the LoRA adapters.
We observed overfitting when training with Full parameter FT: the development loss started deteriorating after a few epochs (e.g., approximately $2$ for SST-2) while the training loss went quickly to zero. We used a combination of early stopping, dropout and weight decay to deal with overfitting. We noticed that SpaRTA is a natural regularizer: increasing sparsity resulted in less overfitting, to a point in which there was no more overfitting (e.g., this was achieved at a sparsity $s>99\%$ for all models under consideration). In general, overfitting was more noticeable with the Mistral 7B models; as expected since they are larger than the Gemma 2B models. We also noticed that larger models require higher sparsity levels to eliminate overfitting given the same training data.
@@ Appendix D Additional Experiment Results
Table 12 confirms targeting Wo and Wv is not the best choice for LoRA (96.2% accuracy for Wo and Wv versus 96.4% for Wq and Wv which is the LoRA default). These results can be compared with the corresponding SpaRTA results in Section 7.2. Tables 13 and Table 14 present the complete set of results on GLUE and BoolQ for the gemma-2b and mistral-7b models, respectively. Tables 15 and 16 present standard errors from corresponding experiments.
LoRA targets rank Loss Accuracy
Wq 16 0.155 94.8 %
Wk 24 0.182 94.5%
Wv 24 0.116 96.2%
Wo 16 0.114 96.2%
Wq, Wk 8 0.152 94.8%
Wq, Wv (default) 8 0.110 96.4%
Wq, Wo 6 0.107 96.6%
Wk, Wv 12 0.110 96.7%
Wk, Wo 8 0.110 96.9%
Wv, Wo 8 0.114 96.2%
Wq, Wv, Wo 4 0.108 96.7%
Wk, Wv, Wo 6 0.113 96.7%
Wq, Wk, Wv, Wo 3 0.108 96.8%
MLP 1 0.112 96.2%
Table 12: Test loss and accuracy on SST-2 after applying LoRA to different types of parameters in the gemma-2b model, given that the same number of trainable parameters are selected. Results averaged across 10 random seeds.
@@
Model: mistral-7b QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ Method Targets loss acc. loss acc. loss acc. loss acc. loss acc. loss acc. loss mcc loss acc. SpaRTA (5%) ALL 0.11 95.9 0.36 85.0 0.11 96.8 0.38 82.1 0.30 89.2 0.33 86.7 0.35 64.6 0.25 90.1 SpaRTA (0.5%) 0.12 95.8 0.44 84.4 0.11 96.7 0.22 90.6 0.26 90.4 0.32 87.0 0.35 66.8 0.26 89.9 SpaRTA (0.048%) 0.13 95.3 0.41 83.6 0.11 96.7 0.23 90.1 0.40 86.7 0.34 86.6 0.36 64.6 0.26 89.8 SpaRTA (0.048%) Q,V 0.14 94.5 0.37 84.0 0.11 96.6 0.24 89.7 0.33 88.6 0.36 85.5 0.36 65.2 0.28 89.3 LoRA (0.048%) 0.12 95.4 0.39 85.4 0.12 96.5 0.23 90.1 0.28 89.7 0.32 87.0 0.38 64.5 0.27 89.6 DoRA (0.048%) 0.12 95.6 0.41 85.3 0.12 96.8 0.24 90.1 0.28 89.7 0.33 86.4 0.34 65.8 0.26 89.9 SpaRTA (0.048%) O,V 0.15 94.5 0.41 85.6 0.21 91.7 0.24 89.6 0.33 87.8 0.35 86.4 0.39 59.8 0.29 89.3
@@
Model: gemma-2b-it QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ Method targets loss acc. loss acc. loss acc. loss acc. loss acc. loss acc. loss mcc loss acc. SpaRTA (5%) ALL 0.00 0.14 0.03 1.54 0.01 0.51 0.00 0.21 0.01 0.26 0.01 0.48 0.00 0.88 0.01 0.54 SpaRTA (0.5%) 0.00 0.16 0.03 0.64 0.01 0.46 0.00 0.17 0.00 0.13 0.01 0.07 0.01 0.75 0.01 0.25 SpaRTA (0.037%) 0.00 0.12 0.01 0.55 0.00 0.17 0.00 0.18 0.00 0.14 0.00 0.45 0.01 0.44 0.00 0.09 LoRA (0.037%) Q,V 0.00 0.11 0.01 1.10 0.00 0.35 0.00 0.03 0.00 0.11 0.00 0.22 0.02 1.40 0.01 0.12 DoRA (0.037%) 0.00 0.11 0.01 0.55 0.00 0.20 0.00 0.03 0.00 0.09 0.00 0.32 0.01 1.31 0.01 0.29 SpaRTA (0.037%) 0.00 0.08 0.00 1.04 0.01 0.11 0.00 0.00 0.00 0.26 0.00 0.22 0.02 2.83 0.00 0.41 SpaRTA (0.037%) O,V 0.00 0.13 0.01 0.64 0.00 0.27 0.00 0.16 0.00 0.17 0.02 0.22 0.01 4.36 0.00 0.22
@@
Model: mistral-7b-it QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ Method targets loss acc. loss acc. loss acc. loss acc. loss acc. loss acc. loss mcc loss acc. SpaRTA (5%) ALL 0.00 0.09 0.01 0.24 0.01 0.26 0.00 0.03 0.00 0.08 0.01 0.47 0.01 1.91 0.00 0.16 SpaRTA (0.5%) 0.00 0.05 0.01 1.07 0.01 0.18 0.00 0.10 0.00 0.09 0.02 0.25 0.01 1.28 0.00 0.09 SpaRTA (0.048%) 0.00 0.10 0.03 0.55 0.00 0.20 0.00 0.03 0.00 0.11 0.01 0.32 0.02 1.87 0.00 0.20 LoRA (0.048%) Q,V 0.00 0.16 0.01 0.60 0.01 0.44 0.00 0.06 0.00 0.13 0.01 0.33 0.01 1.16 0.00 0.17 DoRA (0.048%) 0.02 0.76 0.02 0.87 0.00 0.17 0.00 0.03 0.00 0.07 0.00 0.17 0.01 1.22 0.00 0.16 SpaRTA (0.048%) 0.00 0.06 0.01 0.73 0.00 0.23 0.00 0.04 0.00 0.05 0.01 0.43 0.01 1.17 0.00 0.27 SpaRTA (0.048%) O,V 0.00 0.02 0.01 0.42 0.01 0.34 0.00 0.07 0.00 0.13 0.02 0.24 0.01 0.97 0.00 0.23